# Semantic Context Matters: Improving Conditioning for Autoregressive Models

### v1

@@ 4.2 Controllable Generation Results
C2I Controllable Generation. We evaluate the C2I controllable generation performance of our SCAR and its unified multi-condition variant SCAR-Uni on ImageNet [deng2009imagenet]. As shown in Table 1, we assess both conditional consistency and image generation quality under two autoregressive paradigms: next-token-based LlamaGen and next-set-based VAR. Compared with existing C2I controllable generation methods, including diffusion-based methods [zhang2023adding, mou2024t2i], next-token AR methods [li2024controlar], and next-set AR methods [controlvar, yao2024car], SCAR achieves significantly better image quality (e.g., FID 1.97 vs. 7.69 on Canny), while maintaining competitive control accuracy compared to models that apply decoding-stage injection. SCAR is validated under both next-token and next-set paradigms, demonstrating the effectiveness and generalizability. Figure 3 and Figure 4 further illustrate this, showing visual results of SCAR across different autoregressive paradigms, and SCAR-Uni under a unified multi-condition setting.
Table 3: Quantitative comparison between our SCAR and other methods on PIE-Bench [ju2023direct]. Metrics include structure consistency, background preservation, and CLIP-based image-text alignment. ControlAR* is a retrained variant of ControlAR for editing.
@@
Diff.
InstructPix2Pix [instructpix2pix] SD1.5 107.43 16.69 271.33 392.22 68.39 23.49 22.20
InstructDiffusion [geng2024instructdiffusion] SD1.5 74.21 20.88 142.35 353.45 76.70 24.06 21.57
MGIE [fu2023guiding] SD1.5 67.41 21.20 142.25 295.11 77.52 24.28 21.79
SEED-X-Edit [ge2024seed] SDXL 61.69 18.80 173.63 209.05 75.13 25.51 21.87
AR
ControlAR* [li2024controlar] LlamaGen 116.99 14.63 289.34 590.63 63.03 24.07 23.12
EditAR [mu2025editar] LlamaGen 39.43 21.32 117.15 130.27 75.13 24.87 21.87
SCAR (Ours) LlamaGen 30.98 22.59 105.09 83.47 76.73 26.07 25.08
Figure 6: Qualitative results of instruction editing. Compared with other AR-based methods (ControlAR [li2024controlar], EditAR [mu2025editar]), SCAR can follow instructions more closely and preserve the original image content better. Note: more visualizations can be found in the Appendix.
T2I Controllable Generation. We adopt LlamaGen-XL as AR model for SCAR in T2I controllable generation. Table 2 reports generation quality (FID) and controllability results across four control conditions on MultiGen-20M [unicontrol]. Alongside representative diffusion-based methods [li2023gligen, t2i-adapter, zhang2023adding, uni-controlnet, unicontrol, li2024controlnet++], we compare SCAR with recent AR-based methods, including EditAR [mu2025editar] and ControlAR [li2024controlar]. SCAR consistently outperforms all competing methods in FID across all conditions, achieving notably lower scores (e.g., HED: 8.41 vs. 10.53; Depth: 13.77 vs. 14.61; Our SCAR vs. ControlAR). While the decoding-stage injection method ControlAR achieves strong alignment, SCAR still delivers highly competitive consistency across HED, Depth, and others. Notably, under the same backbone, SCAR achieves a better balance between control accuracy and visual fidelity compared to prior methods. Figure 5 shows qualitative comparisons, where SCAR generates more natural and visually appealing results while maintaining comparable controllability.
@@ 4.3 Instruction Editing Results
Quantitative Results. Table 3 presents quantitative results on PIE-Bench [ju2023direct], evaluating structure preservation, background reconstruction, and image-text alignment. Under the same backbone architecture (LlamaGen-XL), prefilling-based methods (SCAR and EditAR) substantially outperform decoding-stage injection (ControlAR). This indicates decoding-stage injection is suboptimal for instruction editing, as its strong spatial guidance tends to disrupt the autoregressive editing. SCAR achieves the best results on most metrics, reducing structure distance by 21.4%, LPIPS by 10.3%, and MSE by 35.9%, while improving PSNR by 1.27 dB and boosting regional CLIP similarity. These gains stem from SCAR’s compressed semantic prefilling, which replaces VQ tokens with DINO features that better capture structural and semantic cues. Furthermore, Semantic Alignment Guidance further enhances instruction-result consistency.
Qualitative Results. As shown in Figure 6, we qualitatively compare our SCAR with AR-based methods [li2024controlar, mu2025editar]. SCAR consistently follows instructions while better preserving untouched regions, avoiding unnecessary over-editing. Compared to ControlAR and EditAR, it handles fine-grained, structurally complex edits with greater precision, producing clearer boundaries and more natural textures. These results highlight SCAR’s advantages in semantic alignment, structural consistency, and visual quality. More visualizations can be found in the Appendix
@@ 4.4 Ablation Study
MultiGen-20M
Strategy HED Depth
@@
MultiGen-20M
Compression
Ratio $k^{2}$ for Eq. 3
HED Depth
@@
Idx
Image
Encoder $\mathcal{E}$
Para.
of $\mathcal{E}$
HED Depth
@@
Ablations on Image Encoder $\mathcal{E}$. In Table 6, we conduct experiments using different image encoders towards different controls on MultiGen-20M [unicontrol], including HED and Depth. Firstly, unlike previous approaches [li2024controlar, yao2024car], Our methods adopt frozen image encoders [vit, dinov2, clip, sam] to preserve the robust features obtained from large-scale pretraining. As shown in Table 6 (a) to (c), DINOv2 [dinov2] outperforms other pretrained vision models such as ViT [vit], CLIP [clip], and SAM [sam]. Furthermore, comparing (b), (c), (a), and (f) indicates that increasing the scale of the image encoder significantly boosts performance. Therefore, we choose DINOv2-B as the final image encoder.
@@ Appendix A More Discussion
For controllable generation, decoding-stage methods [li2024controlar, yao2024car, xu2025scalar, liu2025scaleweaver] like ControlAR [li2024controlar] and CAR [yao2024car] inject control signals into intermediate layers, enabling fine-grained spatial alignment. However, this tight coupling often leads to training instability and overfitting to local structures, resulting in blurry textures and reduced realism. In contrast, prefilling-based methods such as EditAR [mu2025editar] and our SCAR prepend control features as conditioning tokens before decoding. Though offering slightly weaker pixel-level control, this design leads to sharper, more natural outputs and better generalization across control types. More importantly, for instruction editing, decoding-stage injection fails to align generation with high-level semantics, often producing overconstrained or inconsistent results. It also conflicts with unified multimodal models (UMM) [an2024mc, luo2024llm, ren2024ultrapixel, peng2024efficient, lin2024draw, lin2025perceiveanythingrecognizeexplain, an2025unictokens, xu2025camel], as spatial injection disrupts the shared AR generation. Prefilling-stage conditioning avoids this issue, integrates cleanly with UMM, and better preserves semantic intent during generation.
Semantic Prefix vs. VQ Prefix. Table 7 reports the quantitative comparison between semantically aligned DINO tokens and standard VQ tokens under two compression settings ($k^{2}{=}1\times$ and $4\times$) on MultiGen-20M [unicontrol]. Across both settings, DINO tokens consistently outperform VQ tokens, underscoring the effectiveness of our semantic prefix for autoregressive editing. Without the compression ($1\times$ compression) setting, DINO tokens improve SSIM by +1.96 and reduce FID by 0.91, indicating better perceptual quality and semantic coherence. Under the $4\times$ compression setting, the improvement becomes even more pronounced: SSIM increases by +3.21, and FID drops by 2.83. These results demonstrate that our method not only improves generation quality in standard setups but also maintains strong performance under token compression.
@@
Prefix Condition
Compress
Ratio $k^{2}$
HED
@@
Discussion on Semantic Alignment. Recent work increasingly suggests that enforcing semantic representation alignment can lead to more effective learning in generative models. In diffusion models, REPA [yu2025repa] aligns internal features with pretrained semantic encoders to stabilize training and improve generation. REG [reg] takes a different path by injecting discriminative semantics directly via spatial concatenation with a class token, improving quality and convergence with minimal overhead. VA-VAE [vavae] enhances latent interpretability via representation alignment in VAEs. MaskDiT [zheng2023fast] enforces semantic consistency through masked reconstruction or auxiliary decoders. Multi-stage methods [pernias2023wurstchen, li2024return] leverage pretrained representations as intermediate maps, while USP [chu2025usp] aligns masked latents in a shared VAE space to unify generation and understanding.
While these strategies have shown strong results in diffusion-based models, extending semantic alignment to autoregressive (AR) frameworks for editing presents unique challenges. EditAR [mu2025editar] attempts this by distilling supervision from DINO features onto the hidden states of generated VQ tokens. However, as this supervision is applied post-generation, it lacks explicit semantic flow during decoding.
In contrast, SCAR aligns semantic features at the prefilling stage, by matching conditional representations from the source image with a pretrained visual space (e.g., DINO [dinov2]) before decoding begins. This design fits naturally with the AR model’s causal structure, introducing dense supervision without interfering with token-level predictions. As a result, the model internalizes the semantic correspondence early and propagates it consistently throughout generation. Compared to global alignment strategies in diffusion or post-hoc distillation in EditAR, SCAR provides a more direct and effective mechanism for integrating conditional guidance, especially in instruction-based editing where semantic consistency across text, condition, and output is critical.
Figure 8: Additional visualizations on Semantic Alignment Guidance $\mathcal{L}_{align}$ in Equation 8.
@@ B.3 Controllable Generation
Figure 9: Additional visualizations of instruction editing results. SCAR (Ours) produces more faithful and semantically consistent edits than ControlAR [li2024controlar] and EditAR [mu2025editar], with all methods using the same LlamaGen-XL [llamagen]. All visualizations are generated at a resolution of $\times512$.
@@
Figure 11: Additional visualizations of T2I Controllable Generation by SCAR based on LlamaGen-XL. To demonstrate the controllability and generalization ability of our SCAR, we present results under four control conditions: Canny, Depth, HED, and Lineart. All images are generated at a resolution of $\times512$.

### v2

@@ 4.2 Controllable Generation Results
C2I Controllable Generation. We evaluate the C2I controllable generation performance of our SCAR and its unified multi-condition variant SCAR-Uni on ImageNet [9]. As shown in Table 1, we assess both conditional consistency and image generation quality under two autoregressive paradigms: next-token-based LlamaGen and next-set-based VAR. Compared with existing C2I controllable generation methods, including diffusion-based methods [90, 48], next-token AR methods [32], and next-set AR methods [30, 85], SCAR achieves significantly better image quality (e.g., FID 1.97 vs. 7.69 on Canny), while maintaining competitive control accuracy compared to models that apply decoding-stage injection. SCAR is validated under both next-token and next-set paradigms, demonstrating the effectiveness and generalizability. Figure 3 and Figure 4 further illustrate this, showing visual results of SCAR across different autoregressive paradigms, and SCAR-Uni under a unified multi-condition setting.
Table 3: Quantitative comparison between our SCAR and other methods on PIE-Bench [21]. Metrics include structure consistency, background preservation, and CLIP-based image-text alignment. ControlAR* is a retrained variant of ControlAR for editing.
@@
Diff. InstructPix2Pix [4] SD1.5 107.43 16.69 271.33 392.22 68.39 23.49 22.20
InstructDiffusion [14] SD1.5 74.21 20.88 142.35 353.45 76.70 24.06 21.57
MGIE [12] SD1.5 67.41 21.20 142.25 295.11 77.52 24.28 21.79
SEED-X-Edit [13] SDXL 61.69 18.80 173.63 209.05 75.13 25.51 21.87
AR ControlAR* [32] LlamaGen 116.99 14.63 289.34 590.63 63.03 23.12 21.43
EditAR [50] LlamaGen 39.43 21.32 117.15 130.27 75.13 24.87 21.87
SCAR (Ours) LlamaGen 30.98 22.59 105.09 83.47 76.73 25.08 21.88
Figure 6: Qualitative results of instruction editing. Compared with other AR-based methods (ControlAR [32], EditAR [50]), SCAR can follow instructions more closely and preserve the original image content better. Note: more visualizations can be found in the Appendix.
T2I Controllable Generation. We adopt LlamaGen-XL as AR model for SCAR in T2I controllable generation. Table 2 reports generation quality (FID) and controllability results across four control conditions on MultiGen-20M [58]. Alongside representative diffusion-based methods [31, 49, 90, 94, 58, 28], we compare SCAR with recent AR-based methods, including EditAR [50] and ControlAR [32]. SCAR consistently outperforms all competing methods in FID across all conditions, achieving notably lower scores (e.g., HED: 8.41 vs. 10.53; Depth: 13.77 vs. 14.61; Our SCAR vs. ControlAR). While the decoding-stage injection method ControlAR achieves strong alignment, SCAR still delivers highly competitive consistency across HED, Depth, and others. Notably, under the same backbone, SCAR achieves a better balance between control accuracy and visual fidelity compared to prior methods. Figure 5 shows qualitative comparisons, where SCAR generates more natural and visually appealing results while maintaining comparable controllability.
@@ 4.3 Instruction Editing Results
Quantitative Results. Table 3 presents quantitative results on PIE-Bench [21], evaluating structure preservation, background reconstruction, and image-text alignment. Under the same backbone architecture (LlamaGen-XL), prefilling-based methods (SCAR and EditAR) substantially outperform decoding-stage injection (ControlAR). This indicates decoding-stage injection is suboptimal for instruction editing, as its strong spatial guidance tends to disrupt the autoregressive editing. SCAR achieves the best results on most metrics, reducing structure distance by 21.4%, LPIPS by 10.3%, and MSE by 35.9%, while improving PSNR by 1.27 dB and boosting regional CLIP similarity. These gains stem from SCAR’s compressed semantic prefilling, which replaces VQ tokens with DINO features that better capture structural and semantic cues. Furthermore, Semantic Alignment Guidance further enhances instruction-result consistency.
Qualitative Results. As shown in Figure 6, we qualitatively compare our SCAR with AR-based methods [32, 50]. SCAR consistently follows instructions while better preserving untouched regions, avoiding unnecessary over-editing. Compared to ControlAR and EditAR, it handles fine-grained, structurally complex edits with greater precision, producing clearer boundaries and more natural textures. These results highlight SCAR’s advantages in semantic alignment, structural consistency, and visual quality. More visualizations can be found in the Appendix
@@ 4.4 Ablation Study
MultiGen-20M Strategy HED Depth
@@
MultiGen-20M Compression Ratio $k^{2}$ for Eq. 3 HED Depth
@@
Idx Image Encoder $\mathcal{E}$ Para. of $\mathcal{E}$ HED Depth
@@
Ablations on Image Encoder $\mathcal{E}$. In Table 6, we conduct experiments using different image encoders towards different controls on MultiGen-20M [58], including HED and Depth. Firstly, unlike previous approaches [32, 85], Our methods adopt frozen image encoders [10, 52, 61, 23] to preserve the robust features obtained from large-scale pretraining. As shown in Table 6 (a) to (c), DINOv2 [52] outperforms other pretrained vision models such as ViT [10], CLIP [61], and SAM [23]. Furthermore, comparing (b), (c), (a), and (f) indicates that increasing the scale of the image encoder significantly boosts performance. Therefore, we choose DINOv2-B as the final image encoder.
@@ Appendix A More Discussion
For controllable generation, decoding-stage methods [32, 85, 83, 37] like ControlAR [32] and CAR [85] inject control signals into intermediate layers, enabling fine-grained spatial alignment. However, this tight coupling often leads to training instability and overfitting to local structures, resulting in blurry textures and reduced realism. In contrast, prefilling-based methods such as EditAR [50] and our SCAR prepend control features as conditioning tokens before decoding. Though offering slightly weaker pixel-level control, this design leads to sharper, more natural outputs and better generalization across control types. More importantly, for instruction editing, decoding-stage injection fails to align generation with high-level semantics, often producing overconstrained or inconsistent results. It also conflicts with unified multimodal models (UMM) [2, 38, 63, 54, 33, 34, 3, 82], as spatial injection disrupts the shared AR generation. Prefilling-stage conditioning avoids this issue, integrates cleanly with UMM, and better preserves semantic intent during generation.
Semantic Prefix vs. VQ Prefix. Table 7 reports the quantitative comparison between semantically aligned DINO tokens and standard VQ tokens under two compression settings ($k^{2}{=}1\times$ and $4\times$) on MultiGen-20M [58]. Across both settings, DINO tokens consistently outperform VQ tokens, underscoring the effectiveness of our semantic prefix for autoregressive editing. Without the compression ($1\times$ compression) setting, DINO tokens improve SSIM by +1.96 and reduce FID by 0.91, indicating better perceptual quality and semantic coherence. Under the $4\times$ compression setting, the improvement becomes even more pronounced: SSIM increases by +3.21, and FID drops by 2.83. These results demonstrate that our method not only improves generation quality in standard setups but also maintains strong performance under token compression.
@@
Prefix Condition Compress Ratio $k^{2}$ HED
@@
Discussion on Semantic Alignment. Recent work increasingly suggests that enforcing semantic representation alignment can lead to more effective learning in generative models. In diffusion models, REPA [86] aligns internal features with pretrained semantic encoders to stabilize training and improve generation. REG [79] takes a different path by injecting discriminative semantics directly via spatial concatenation with a class token, improving quality and convergence with minimal overhead. VA-VAE [84] enhances latent interpretability via representation alignment in VAEs. MaskDiT [95] enforces semantic consistency through masked reconstruction or auxiliary decoders. Multi-stage methods [55, 29] leverage pretrained representations as intermediate maps, while USP [8] aligns masked latents in a shared VAE space to unify generation and understanding.
While these strategies have shown strong results in diffusion-based models, extending semantic alignment to autoregressive (AR) frameworks for editing presents unique challenges. EditAR [50] attempts this by distilling supervision from DINO features onto the hidden states of generated VQ tokens. However, as this supervision is applied post-generation, it lacks explicit semantic flow during decoding.
In contrast, SCAR aligns semantic features at the prefilling stage, by matching conditional representations from the source image with a pretrained visual space (e.g., DINO [52]) before decoding begins. This design fits naturally with the AR model’s causal structure, introducing dense supervision without interfering with token-level predictions. As a result, the model internalizes the semantic correspondence early and propagates it consistently throughout generation. Compared to global alignment strategies in diffusion or post-hoc distillation in EditAR, SCAR provides a more direct and effective mechanism for integrating conditional guidance, especially in instruction-based editing where semantic consistency across text, condition, and output is critical.
Figure 8: Additional visualizations on Semantic Alignment Guidance $\mathcal{L}_{align}$ in Equation (8).
@@ B.3 Controllable Generation
Figure 9: Additional visualizations of instruction editing results. SCAR (Ours) produces more faithful and semantically consistent edits than ControlAR [32] and EditAR [50], with all methods using the same LlamaGen-XL [69]. All visualizations are generated at a resolution of $\times512$.
@@
Figure 11: Additional visualizations of T2I Controllable Generation by SCAR based on LlamaGen-XL. To demonstrate the controllability and generalization ability of our SCAR, we present results under four control conditions: Canny, Depth, HED, and Lineart. All images are generated at a resolution of 512×512.